### v2

# Parallel methods for quasinonexpansive mappings in a Hilbert space

Abstract.
This paper is devoted to the problem of finding a common fixed point of quasinonexpansive mappings defined on a Hilbert space. To approximate the solution to this problem, we present several iterative processes using the parallel method based on [MR3196179, JMM2018].
Key words and phrases:
Parallel method, quasinonexpansive mapping, fixed point, approximation algorithm
2010 Mathematics Subject Classification:
47J25, 47J20, 47H09
1. Introduction
The objective of this study is to solve the following common fixed point problem:
Find $z\in\bigcap_{i=1}^{N}\operatorname{F}(T_{i})$,
where $N$ is a positive integer, $T_{i}$ is a quasinonexpansive mapping of a subset $C$ of a Hilbert space $H$, and $\operatorname{F}(T_{i})$ is the set of fixed points of $T_{i}$ for $i\in\{1,\dotsc,N\}$. In particular, we intend to prove that if $u$ is a given point in $H$, then the sequence $\{x_{n}\}$ generated by $u$ and $T_{i}$ converges strongly to the common fixed point of $\{T_{i}\}$ closest to $u$.
In the previous study [JMM2018], we established some strong convergence results regarding the problem by the parallel hybrid method and the parallel shrinking method; see [MR3196179] and the references therein for the parallel hybrid method. These results were proven using convergence theorems obtained in [MR2884574] and some properties [JMM2018]*Lemmas 2.2 and 2.3 of the parallel algorithm.
In the present study, we obtain other convergence results for the common fixed point problem by using the parallel method based on [MR3196179, JMM2018] and some simple iteration processes developed in [MR2529497, MR2960628].
The paper is organized as follows: Section 2 introduces some relevant notions, definitions, and lemmas. In Section 3, we prove a strong convergence theorem for the common fixed point problem of quasinonexpansive nonself-mappings. In Section 4, we also provide strong and weak convergence theorems for the common fixed point problem of strongly quasinonexpansive mappings.
2. Preliminaries
Throughout the present paper, $H$ denotes a real Hilbert space, $\left\langle\,\cdot\,$$\cdot\,\right\rangle$ the inner product of $H$, $\left\lVert\,\cdot\,\right\rVert$ the norm of $H$, $C$ a nonempty subset of $H$, II the identity mapping on $H$, and $\mathbb{N}$ the set of positive integers. Strong convergence of a sequence $\{x_{n}\}$ in $H$ to $z\in H$ is denoted by $x_{n}\to z$ and weak convergence by $x_{n}\rightharpoonup z$.
Let $T\colon C\to H$ be a mapping. The set of fixed points of $T$ is denoted by $\operatorname{F}(T)$, that is, $\operatorname{F}(T)=\{z\in C\colon Tz=z\}$. A point $z\in C$ is said to be an asymptotic fixed point of $T$ [MR2058234] if there exists a sequence $\{x_{n}\}$ in $C$ such that $x_{n}-Tx_{n}\to 0$ and $x_{n}\rightharpoonup z$. The set of asymptotic fixed points of $T$ is denoted by $\hat{\mathrm{F}}(T)$. It is clear that $\operatorname{F}(T)\subset\hat{\mathrm{F}}(T)$. A mapping $T$ is said to be quasinonexpansive if $\operatorname{F}(T)\neq\emptyset$ and $\left\lVert Tx-z\right\rVert\leq\left\lVert x-z\right\rVert$ for all $x\in C$ and $z\in\operatorname{F}(T)$; $T$ is said to be nonexpansive if $\left\lVert Tx-Ty\right\rVert\leq\left\lVert x-y\right\rVert$ for all $x,y\in C$; $T$ is said to be firmly nonexpansive [MR0215141] if $\left\langle x-y,Tx-Ty\right\rangle\geq\left\lVert Tx-Ty\right\rVert^{2}$, or equivalently
(2.1) $\left\lVert Tx-Ty\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert x-y-(Tx-Ty)\right\rVert^{2}$
for all $x,y\in C$; $T$ is said to be strongly quasinonexpansive [pNACA2017, pNACA2015, MR4136436, MR3908455, MR3213161] if $T$ is quasinonexpansive and $Tx_{n}-x_{n}\to 0$ whenever $\{x_{n}\}$ is a bounded sequence in $C$ and $\left\lVert x_{n}-p\right\rVert-\left\lVert Tx_{n}-p\right\rVert\to 0$ for some point $p\in\operatorname{F}(T)$; $T$ is said to be demiclosed at 0 if $z\in C$ and $Tz=0$ whenever $\{x_{n}\}$ is a sequence in $C$ such that $x_{n}\rightharpoonup z$ and $Tx_{n}\to 0$; see, for example, [MR1074005]. It is clear that ITI-T is demiclosed at 0 if and only if $\hat{\mathrm{F}}(T)=\operatorname{F}(T)$. Moreover, under the assumption that $C$ is closed and convex, we know the following:
  • If $T$ is quasinonexpansive, then $\operatorname{F}(T)$ is closed and convex; see [MR0298499]*Theorem 1;
  • if $T$ is nonexpansive, then $\hat{\mathrm{F}}(T)=\operatorname{F}(T)$; see [MR1074005].
Remark 2.1.
The notion of an asymptotic fixed point above originates from that proposed by Reich [MR1386686]. However, the two notions are slightly different.
Remark 2.2.
The notion of a strongly quasinonexpansive mapping is based on that of a strongly nonexpansive mapping in the sense of [MR1386686]; see also [MR0470761].
Remark 2.3.
If a strongly nonexpansive mapping in the sense of [MR0470761, MR2581778, MR2799767, MR2671943, MR2377867] has a fixed point, then it is strongly quasinonexpansive above.
Remark 2.4.
Let $T\colon C\to H$ be a mapping with a fixed point. Since $\beta_{n}-\alpha_{n}\to 0\Leftrightarrow\beta_{n}^{2}-\alpha_{n}^{2}\to 0$ for all bounded sequences $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ in $0,\infty)$, $T$ is strongly quasinonexpansive if and only if $T$ is of type (sr) in the sense of [MR3258665, MR2529497, MR2884574]; see also [MR2358984, MR3745220].
Remark 2.5.
It is known that the subgradient projection in the sense of [MR3213161] is a strongly quasinonexpansive mapping which is not nonexpansive; see also [MR1895827, MR3745220].
Let $D$ be a nonempty closed convex subset of $H$. It is known that for each $x\in H$ there exists a unique point $x_{0}\in D$ such that
$\left\lVert x-x_{0}\right\rVert=\min\{\left\lVert x-y\right\rVert:y\in D\}.$
Such a point $x_{0}$ is denoted by $P_{D}(x)$ and $P_{D}$ is called the metric projection of $H$ onto $D$. It is known that $P_{D}$ is firmly nonexpansive; see, for example, [MR2548424].
Remark 2.6.
Let $D$ be a nonempty closed convex subset of $H$. Since a firmly nonexpansive mapping is nonexpansive by (2.1), it follows that the metric projection $P_{D}$ of $H$ onto $D$ is nonexpansive. Thus, $\hat{\mathrm{F}}(P_{D})=\operatorname{F}(P_{D})=D$. Moreover, (2.1) also shows that a firmly nonexpansive mapping with a fixed point is strongly quasinonexpansive. Therefore, $P_{D}$ is strongly quasinonexpansive.
Let $\{T_{n}\}$ be a sequence of mappings of $C$ into $H$ such that $\bigcap_{n}\operatorname{F}(T_{n})$ is nonempty. Then
  • $\{T_{n}\}$ is said to be a strongly quasinonexpansive sequence if each $T_{n}$ is quasinonexpansive and $T_{n}x_{n}-x_{n}\to 0$ whenever $\{x_{n}\}$ is a bounded sequence in $C$ and $\left\lVert x_{n}-p\right\rVert-\left\lVert T_{n}x_{n}-p\right\rVert\to 0$ for some point $p\in\bigcap_{n}\operatorname{F}(T_{n})$;
  • a point $z\in C$ is said to be an asymptotic fixed point of $\{T_{n}\}$ if there exist a sequence $\{x_{n}\}$ in $C$ and a subsequence $\{x_{n_{i}}\}$ of $\{x_{n}\}$ such that $T_{n}x_{n}-x_{n}\to 0$ and $x_{n_{i}}\rightharpoonup z$; see [MR3698250, pNACA2017, MR3013135].
The set of asymptotic fixed points of $\{T_{n}\}$ is denoted by $\hat{\operatorname{F}}(\{T_{n}\})$. It is clear that $\bigcap_{n}\operatorname{F}(T_{n})\subset\hat{\operatorname{F}}(\{T_{n}\})$.
Remark 2.7.
We know that a strongly quasinonexpansive sequence above is a sequence of strongly quasinonexpansive type in the sense of [MR3698250, MR3213161, pNACA2017] and a strongly relatively nonexpansive sequence in the sense of [MR3258665, MR2960628, MR2529497, MR3013135]; see [MR3213161]*Remark 2.5. We also know that if a strongly nonexpansive sequence in the sense of [MR2377867, MR2581778, MR2799767] has a common fixed point, then it is a strongly quasinonexpansive sequence above; see also [MR3714903].
Remark 2.8.
We know that the condition $\hat{\mathrm{F}}(\{T_{n}\})=\bigcap_{n}\operatorname{F}(T_{n})$ is equivalent to the condition (Z) in the sense of [MR3258665, MR3185784, MR3213161, MR2960628, MR2799767, MR2671943, MR2529497, MR3203624, pNACA2011, pNLMUA2011, MR2762191, MR2762173, MR3013135], the condition (Z1) in the sense of [pNAO-Asia2008], and the condition (A) in the sense of [MR2581778, MR2884574]; see [MR3013135]*Proposition 6.
Remark 2.9.
We can find some examples of strongly quasinonexpansive sequences with the condition (Z) in [MR3698250], [MR3213161]*Example 4.5, and [MR2799767, MR2529497, MR2960628]; see also [MR2762191, MR2762173, MR3203624, pNACA2011, MR2671943, MR2780284, pNAO-Asia2008, MR2581778, MR2377867, MR3258665, MR3013135].
The following lemma is a direct consequence of [MR2529497]*Theorem 3.8 and  [MR3013135]*Proposition 4; see also [MR2671943]*Lemma 2.2.
Lemma 2.10.
Let $H$ be a Hilbert space, $C$ a nonempty subset of $H$, $\{U_{n}\}$ be a sequence of mappings of $C$ into $H$ such that $\bigcap_{n}\operatorname{F}(U_{n})$ is nonempty, $\{\beta_{n}\}$ a sequence in $(0,1)$, and $S_{n}\colon C\to H$ a mapping defined by $S_{n}=\beta_{n}I+(1-\beta_{n})U_{n}$ for $n\in\mathbb{N}$. Then $\operatorname{F}(U_{n})=\operatorname{F}(S_{n})$ for all $n\in\mathbb{N}$, and moreover, the following hold:
  • If each $U_{n}$ is quasinonexpansive and $\inf_{n}\beta_{n}>0$, then $\{S_{n}\}$ is a strongly quasinonexpansive sequence;
  • if $\sup_{n}\beta_{n}<1$ and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$, then $\hat{\mathrm{F}}(\{S_{n}\})=\bigcap_{n}\operatorname{F}(S_{n})$.
The following lemma is important in the convergence analysis of the parallel method in Sections 3 and 4.
Lemma 2.11.
Let $H$ be a Hilbert space, $C$ a nonempty subset of $H$, $\{U_{n}\}$ be a sequence of mappings of $C$ into $H$, NN a positive integer, and $T_{i}$ a mapping of $C$ into $H$ for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Suppose that $\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$ is nonempty, and that for any $x\in C$ and $n\in\mathbb{N}$ there exists $k\in\arg\max\{\left\lVert T_{i}x-x\right\rVert\colon i\in\Lambda\}$ such that $U_{n}x=T_{k}x$. Then the following hold:
  1. (1)
    $\operatorname{F}(U_{n})=\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$ for all $n\in\mathbb{N}$, and hence $\bigcap_{n}\operatorname{F}(U_{n})$ is nonempty;
  2. (2)
    if $T_{i}$ is quasinonexpansive for all $i\in\Lambda$, then so is $U_{n}$ for all $n\in\mathbb{N}$;
  3. (3)
    if $\hat{\mathrm{F}}(T_{i})=\operatorname{F}(T_{i})$ for all $i\in\Lambda$, then $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$;
  4. (4)
    if $T_{i}$ is a strongly quasinonexpansive mapping for all $i\in\Lambda$, then $\{U_{n}\}$ is a strongly quasinonexpansive sequence.
Proof.
Because (1), (2), and (3) are direct consequences of [JMM2018]*Lemmas 2.2 and 2.3, we must only prove (4). From (1) and (2), we know that $\bigcap_{n}\operatorname{F}(U_{n})=\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$ is nonempty and each $U_{n}$ is quasinonexpansive. Let $\{x_{n}\}$ be a bounded sequence in $C$ and suppose that $\left\lVert x_{n}-p\right\rVert-\left\lVert U_{n}x_{n}-p\right\rVert\to 0$ for some $p\in\bigcap_{n}\operatorname{F}(U_{n})$. We denote $U_{n}x_{n}-x_{n}$ by $z_{n}$. It is enough to show that $z_{n}\to 0$. Let $\{z_{n_{m}}\}$ be a subsequence of $\{z_{n}\}$ and $\{U_{n_{m}}x_{n_{m}}\}$ a subsequence of $\{U_{n}x_{n}\}$ with the same index. By assumption, it turns out that there exist $i\in\Lambda$ and a subsequence $\{U_{n_{m_{l}}}x_{n_{m_{l}}}\}$ of $\{U_{n_{m}}x_{n_{m}}\}$ such that $U_{n_{m_{l}}}x_{n_{m_{l}}}=T_{i}x_{n_{m_{l}}}$ for all $l\in\mathbb{N}$. Since $p\in\operatorname{F}(T_{i})$, $T_{i}$ is strongly quasinonexpansive, and
$\left\lVert x_{n_{m_{l}}}-p\right\rVert-\left\lVert T_{i}x_{n_{m_{l}}}-p\right\rVert=\left\lVert x_{n_{m_{l}}}-p\right\rVert-\left\lVert U_{n_{m_{l}}}x_{n_{m_{l}}}-p\right\rVert\to 0$
as $l\to\infty$, it follows that $z_{n_{m_{l}}}=T_{i}x_{n_{m_{l}}}-x_{n_{m_{l}}}\to 0$. This completes the proof. ∎
The following theorem is a direct consequence of [MR2960628]*Theorem 4.1; see also [pNACA2017, pNACA2011, MR3213161, MR3698250].
Theorem 2.12.
Let $H$ be a Hilbert space, $C$ a nonempty closed convex subset of $H$, $u$ a point in $H$, $\{\alpha_{n}\}$ a sequence in $(0,1]$, $\{S_{n}\}$ a sequence of mappings of $C$ into $H$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
$x_{n+1}=P_{C}\bigl{(}\alpha_{n}u+(1-\alpha_{n})S_{n}x_{n}\bigr{)}$
for $n\in\mathbb{N}$. Suppose that $F=\bigcap_{n}\operatorname{F}(S_{n})\neq\emptyset$, $\{S_{n}\}$ is a strongly quasinonexpansive sequence, $\hat{\operatorname{F}}(\{S_{n}\})=F$, $\sum_{n}\alpha_{n}=\infty$, and $\lim_{n}\alpha_{n}=0$. Then $\{x_{n}\}$ converges strongly to $P_{F}(u)$.
The following theorem is a direct consequence of [MR2529497]*Theorem 4.1; see also [MR2377867].
Theorem 2.13.
Let $H$ be a Hilbert space, $C$ a nonempty closed convex subset of $H$, $\{S_{n}\}$ a sequence of mappings of $C$ into itself, and $\{x_{n}\}$ a sequence in $C$ defined by $x_{1}\in C$ and $x_{n+1}=S_{n}x_{n}$ for $n\in\mathbb{N}$. Suppose that $F=\bigcap_{n}\operatorname{F}(S_{n})\neq\emptyset$, $\{S_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\operatorname{F}}(\{S_{n}\})=F$. Then $\{x_{n}\}$ converges weakly to the strong limit of $\{P_{F}(x_{n})\}$.
3. Parallel methods for quasinonexpansive mappings
In this section, using the parallel method and Theorem 2.12, we prove a strong convergence theorem for the common fixed point problem of quasinonexpansive nonself-mappings.
Theorem 3.1.
Let $H$ be a Hilbert space, $C$ a nonempty closed convex subset of $H$, NN a positive integer, and $T_{i}\colon C\to H$ a quasinonexpansive mapping for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Let $F$ be the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, that is, $F=\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
$\begin{cases}i_{n}\in\arg\max\{\left\lVert T_{i}x_{n}-x_{n}\right\rVert\colon i\in\Lambda\};\\ x_{n+1}=P_{C}\Bigl{(}\alpha_{n}u+(1-\alpha_{n})\bigl{(}\beta_{n}x_{n}+(1-\beta_{n})T_{i_{n}}x_{n}\bigr{)}\Bigr{)}\end{cases}$
for $n\in\mathbb{N}$, where $u\in H$, $\{\alpha_{n}\}$ is a sequence in $(0,1]$, and $\{\beta_{n}\}$ is a sequence in $(0,1)$. Suppose that $F$ is nonempty, $\hat{\mathrm{F}}(T_{i})=\operatorname{F}(T_{i})$ for all $i\in\Lambda$, $\sum_{n}\alpha_{n}=\infty$, $\lim_{n}\alpha_{n}=0$, $\inf_{n}\beta_{n}>0$, and $\sup_{n}\beta_{n}<1$. Then $\{x_{n}\}$ converges strongly to $P_{F}(u)$.
Proof.
Let $M$ be a set-valued mapping of $C$ into $\Lambda$ defined by
$M(x)=\arg\max\{\left\lVert T_{i}x-x\right\rVert\colon i\in\Lambda\}$
for $x\in C$ and let $U_{n}\colon C\to H$ be a mapping defined by
$U_{n}x=\begin{cases}T_{i_{n}}x&\text{if $x=x_{n}$;}\\ T_{j(x)}x&\text{otherwise}\end{cases}$
for $x\in C$ and $n\in\mathbb{N}$, where $j(x)=\min M(x)$ for $x\in C$. Then we can verify that for any $x\in C$ and $n\in\mathbb{N}$, there exists $k\in M(x)$ such that $U_{n}x=T_{k}x$. Thus Lemma 2.11 implies that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, each $U_{n}$ is quasinonexpansive, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. Set $S_{n}=\beta_{n}I+(1-\beta_{n})U_{n}$ for $n\in\mathbb{N}$. Then Lemma 2.10 shows that $\{S_{n}\}$ is a strongly quasinonexpansive sequence and $\hat{\mathrm{F}}(\{S_{n}\})=\bigcap_{n}\operatorname{F}(S_{n})=F$. From the definition of $U_{n}$, it is clear that $x_{n+1}=P_{C}\bigl{(}\alpha_{n}u+(1-\alpha_{n})S_{n}x_{n}\bigr{)}$ for all $n\in\mathbb{N}$. Using Theorem 2.12, we conclude that $\{x_{n}\}$ converges strongly to $P_{F}(u)$. ∎
If we suppose that every $T_{i}$ is a self-mapping and $u\in C$ in Theorem 3.1, then we deduce the following:
Corollary 3.2.
Let $H$, $C$, NN, $\Lambda$, $\{\alpha_{n}\}$, and $\{\beta_{n}\}$ be the same as in Theorem 3.1. Let $T_{i}\colon C\to C$ be a quasinonexpansive mapping for $i\in\Lambda$, $F$ the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
$\begin{cases}i_{n}\in\arg\max\{\left\lVert T_{i}x_{n}-x_{n}\right\rVert\colon i\in\Lambda\};\\ x_{n+1}=\alpha_{n}u+(1-\alpha_{n})\bigl{(}\beta_{n}x_{n}+(1-\beta_{n})T_{i_{n}}x_{n}\bigr{)}\end{cases}$
for $n\in\mathbb{N}$, where $u\in C$. Suppose that $F$ is nonempty and $\hat{\mathrm{F}}(T_{i})=\operatorname{F}(T_{i})$ for all $i\in\Lambda$. Then $\{x_{n}\}$ converges strongly to $P_{F}(u)$.
4. Parallel methods for strongly quasinonexpansive mappings
In this section, we prove strong and weak convergence theorems for the common fixed point problem of strongly quasinonexpansive self-mappings. We then explain how the theorems can be applied to common fixed point problems for strongly quasinonexpansive nonself-mappings and quasinonexpansive mappings.
Using Lemma 2.11 and Theorem 2.12, we obtain the following:
Theorem 4.1.
Let $H$ be a Hilbert space, $N$ a positive integer, $S_{i}\colon H\to H$ a strongly quasinonexpansive mapping for $i\in\Lambda$, and $F$ the set of common fixed points of $\{S_{i}\}_{i\in\Lambda}$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Suppose that $F$ is nonempty and $\hat{\mathrm{F}}(S_{i})=\operatorname{F}(S_{i})$ for all $i\in\Lambda$. Let $\{x_{n}\}$ be a sequence defined by $x_{1}\in H$ and
$\begin{cases}i_{n}\in\arg\max\{\left\lVert S_{i}x_{n}-x_{n}\right\rVert\colon i\in\Lambda\};\\ x_{n+1}=\alpha_{n}u+(1-\alpha_{n})S_{i_{n}}x_{n}\end{cases}$
for $n\in\mathbb{N}$, where $u\in H$ and $\{\alpha_{n}\}$ is a sequence in $(0,1]$. If $\sum_{n}\alpha_{n}=\infty$ and $\lim_{n}\alpha_{n}=0$, then $\{x_{n}\}$ converges strongly to $P_{F}(u)$.
Proof.
Let $M$ be a set-valued mapping of $H$ into $\Lambda$ defined by
(4.1) $M(x)=\arg\max\{\left\lVert S_{i}x-x\right\rVert\colon i\in\Lambda\}$
for $x\in H$ and let $U_{n}\colon H\to H$ be a mapping defined by
(4.2) $U_{n}x=\begin{cases}S_{i_{n}}x&\text{if $x=x_{n}$;}\\ S_{j(x)}x&\text{otherwise}\end{cases}$
for $x\in H$ and $n\in\mathbb{N}$, where $j(x)=\min M(x)$ for $x\in H$. Then we can verify that for any $x\in H$ and $n\in\mathbb{N}$, there exists $k\in M(x)$ such that $U_{n}x=S_{k}x$. Thus Lemma 2.11 implies that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, $\{U_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. From the definition of $U_{n}$, it is clear that $x_{n+1}=\alpha_{n}u+(1-\alpha_{n})U_{n}x_{n}$ for all $n\in\mathbb{N}$. Using Theorem 2.12, we conclude that $\{x_{n}\}$ converges strongly to $P_{F}(u)$. ∎
Using Lemma 2.11 and Theorem 2.13, we obtain the following weak convergence result:
Theorem 4.2.
Let $H$, NN, $S_{i}$, $\Lambda$, $F$ be the same as in Theorem 4.1 and $\{x_{n}\}$ a sequence defined by $x_{1}\in H$ and
$\begin{cases}i_{n}\in\arg\max\{\left\lVert S_{i}x_{n}-x_{n}\right\rVert\colon i\in\Lambda\};\\ x_{n+1}=S_{i_{n}}x_{n}\end{cases}$
for $n\in\mathbb{N}$. Then $\{x_{n}\}$ converges weakly to the strong limit of $\{P_{F}(x_{n})\}$.
Proof.
Let $M$ be a set-valued mapping of $H$ into $\Lambda$ defined by (4.1) for $x\in H$ and $U_{n}\colon H\to H$ a mapping defined by (4.2) for $x\in H$ and $n\in\mathbb{N}$. As in the proof of Theorem 4.1, it follows from Lemma 2.11 that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, $\{U_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. From the definition of $U_{n}$, it is clear that $x_{n+1}=U_{n}x_{n}$ for all $n\in\mathbb{N}$. Applying Theorem 2.13, we complete the proof. ∎
Remark 4.3.
Theorems 4.1 and 4.2 are applicable to the construction of a common fixed point of a finite family of strongly quasinonexpansive nonself-mappings. Indeed, let $C$ be a nonempty closed convex subset of a Hilbert space $H$ and $S\colon C\to H$ a strongly quasinonexpansive mapping. Then it is clear that $SP_{C}$ is a mapping of $H$ into $H$. Furthermore, it is known that $\operatorname{F}(SP_{C})=\operatorname{F}(S)$ and $SP_{C}$ is strongly quasinonexpansive; see [MR2884574]*Lemmas 3.2 and 3.3 and  [MR3908455]*Theorem 5.5. Moreover, it is known that $\hat{\mathrm{F}}(SP_{C})=\operatorname{F}(SP_{C})$ if $\hat{\mathrm{F}}(S)=\operatorname{F}(S)$; see [MR2884574]*Theorem 3.4 and Remark 2.6.
Remark 4.4.
Theorems 4.1 and 4.2 are applicable to the construction of a common fixed point of a finite family of quasinonexpansive mappings. Indeed, let $T\colon H\to H$ be a quasinonexpansive mapping. Set $S_{\lambda}=\lambda I+(1-\lambda)T$ for some $\lambda\in(0,1)$. Then it is known that $\operatorname{F}(S_{\lambda})=\operatorname{F}(T)$ and $S_{\lambda}$ is strongly quasinonexpansive; see [MR2358984]*Corollary 3.4. Moreover, it is known that $\hat{\mathrm{F}}(S_{\lambda})=\operatorname{F}(S_{\lambda})$ if $\hat{\mathrm{F}}(T)=\operatorname{F}(T)$; see [MR2884574]*Corollary 3.8.